Stephaney Company produces several products in its factory, including a karate robe. The company uses a standard cost system to assist in the control of costs. According to the standards that have been set for the robes, the factory should work 780 direct labor-hours each month and produce 2,600 robes. The standard costs associated with this level of production are as follows:

Total Per Unit
of Product
  Direct materials $ 66,300    $ 25.50   
  Direct labor $ 10,920    4.20   
  Variable manufacturing overhead
  (based on direct labor-hours)
$ 1,560    0.60   
$ 30.30   

  

    During April, the factory worked only 720 direct labor-hours and produced 2,700 robes. The following actual costs were recorded during the month:

  

Total Per Unit
of Product
  Direct materials (10,800 yards) $ 68,040    $ 25.20   
  Direct labor $ 11,880    4.40   
  Variable manufacturing overhead $ 5,670    2.10   
$ 31.70   

  

At standard, each robe should require 3.4 yards of material. All of the materials purchased during the month were used in production.

Compute the materials price and quantity variances for April: (Input all amounts as positive values. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Round your intermediate calculations to 2 decimal places. Round your final answers to the nearest dollar amount.)

Materials price variance   
Materials quantity variance

Compute the labor rate and efficiency variances for April: (Input all amounts as positive values. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Round your intermediate calculations to 2 decimal places. Round your final answers to the nearest dollar amount.)

Labor rate variance
Labor efficiency variance

Compute the variable manufacturing overhead rate and efficiency variances for April: (Input all amounts as positive values. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Do not round intermediate calculations. Round your final answers to the nearest dollar amount.)

Variable overhead rate variance
Variable overhead efficiency variance
Answer #1

1)

material price variance  

(AP -SP ) AQ

actual price = 6.3   ( 68040 / 10800 )

standard price = 7.5   (25.5 / 3.4 )

actual quantity = 10800

(7.5 - 6.3 ) * 10800 = 12960

material price variance = 12960 it is favorable (AP < SP )

2)

MATERIAL QUANTITY VARIANCE

(AQ - SQ ) SP

actual quantity = 10800

standard quantity =9990 (2700 * 3.4 )

standard price = 7.5

(10800 - 9990 ) 7.5 = 6075

material quantity variance = 6075 it is unfavorable ( AQ> SQ )

3)

LABOR RATE VARIANCE

(AR - SR ) * AH

actual rate = 16.5 ( 11880 / 720 )

standard rate = 14 ( 10920 / 780 )

actual hour = 720

(16.5 - 14 ) * 720 = 1800

labor rate variance = 1800 it is unfavorable ( AR > SR )

4)

LABOR EFFICIENCY VARIANCE

(AH - SH ) * SR

actual hour = 720

standard hour = 810 (it is flexible ) (780 / 2600 = 0.3 , 03 * 2700 )

standard rate = 14

( 810 - 720 ) * 14 = 1260

efficiency variance = 1260 it is favorable ( AH < SH )

5)

VARIABLE OVERHEAD PRICE VARIANCE

( AR - SR ) AH

actual rate = 7.875   (5670 / 720 )

standard rate = 2   (1560 / 780 )

actual hour = 720

( 7.875 - 2 ) * 720 = 4230

variable overhead rate variance = 4230 it is unfavorable ( AR> SR )

6) VARIABLE OVERHEAD EFFICIENCY VARIANCE

( AH - SH )SR

actual hour =720

standard hour = 810 (780 / 2600 = 0.3 ) ,( 0.3 * 2700=810)

standard rate = 2

( 810 - 720 ) * 2 = 180

variable overhead efficiency variance = 180 it is favorable ( AH < SH )